### v1

@@ 4.2 Limit Posteriors
The second claim in Lemma 4 presents this restriction: If agent $L$ updates down on a state where agent $H$ updates up, then it cannot also be the case that $L$ updates up on a state on a state where $H$ updates down. This observation is also implicitly used by BHK. The first claim is of independent interest, as it conveys the key message of this section in the context of general (unordered) state spaces: Bayesian updating imposes only mild restrictions on belief polarization, allowing beliefs to diverge on all but two states. [^16] Note that updating in opposite directions is not synonymous with polarization as agents may update in opposite directions towards each other. But it is straightforward to construct an example where agents’ beliefs move away from each other on all but two states.
@@
Let $P,P^{\prime}\in\Delta(\Theta)$ be full support priors over a finite state space $\Theta$ and $Q,Q^{\prime}$ be the corresponding posteriors following signal realization $x$. If beliefs move at all then:
  1. 1.
    There exist two states—the minimum and maximum likelihood states—on which beliefs move in the same direction. On all other states, beliefs may move in opposite directions.
  2. 2.
    If beliefs move in opposite directions, the divergence is consistent across states: If there exists a state $\theta\in\Theta$ for which $Q(\theta)-P(\theta)\leq 0\leq Q^{\prime}(\theta)-P^{\prime}(\theta)$, with one of these inequalities strict, then there does not exist a state $\theta^{\prime}$ at which $Q(\theta^{\prime})-P(\theta^{\prime})\geq 0\geq Q^{\prime}(\theta^{\prime})-P^{\prime}(\theta^{\prime})$.
Proof. Recall that $\ell(\theta):=\Pr[X=x|\theta]$ is the likelihood function. Let $\theta^{-}\in\arg\min(\ell)$ and $\theta^{+}\in\arg\max(\ell)$ be a min- and a max-likelihood state. Let $E_{P}[\ell],E_{P^{\prime}}[\ell]$ denote the expectations of $\ell$ according to $P$ and $P^{\prime}$, respectively. By Lemma 9 in the appendix, for every $\theta\in\Theta$ we have $Q(\theta)>P(\theta)$ if and only if $\ell(\theta)>E_{P}[\ell]$, and similarly for $Q^{\prime},P^{\prime}$. For beliefs to move at all, the likelihood function must not be constant, which implies $\ell(\theta^{-})<\ell(\theta^{+})$. Thus, since $P$ and $P^{\prime}$ have full support, $\ell(\theta^{-})<\min\{E_{P}[\ell],E_{P^{\prime}}[\ell]\}$ and $\ell(\theta^{+})>\max\{E_{P}[\ell],E_{P^{\prime}}[\ell]\}$. So beliefs move in the same direction on $\theta^{-},\theta^{+}$.
Finally, assume that there exists $\theta\in\Theta$ for which $Q(\theta)-P(\theta)<0\leq Q^{\prime}(\theta)-P^{\prime}(\theta)$ (here the first inequality is the strict one w.l.o.g). Then, $E_{P}[\ell]>\ell(\theta)\geq E_{P^{\prime}}[\ell]$. So, there does not exist $\theta^{\prime}\in\Theta$ at which $Q(\theta^{\prime})-P(\theta^{\prime})\geq 0\geq Q^{\prime}(\theta^{\prime})-P^{\prime}(\theta^{\prime})$, as that would imply $E_{P}[\ell]\leq E_{P^{\prime}}[\ell]$, a contradiction.
Lemma 4 highlights two restrictions on belief divergence in the face of public information. Yet agents can still polarize through two channels. The first is state-level polarization: beliefs can diverge on all but two states. The second is event-level polarization: even when beliefs move in the same directions on every state, they may diverge on events because the magnitude of belief updates depends on the priors. For one agent, small decreases on some states within an event may be outweighed by large increases on others. For the other, the opposite may hold, leading to polarization on the event.
In the long run, the first channel disappears: for any two full-support priors, limit posteriors move in the same direction on every state, placing positive mass on the identified set and zero elsewhere. Hence, limit coordinatewise polarization operates through the second channel. By contrast, upper orthant and stochastic polarization require divergence on at least one individual state. This is impossible for limit beliefs, but in the short run the combination of the first and second channels allows for upper orthant polarization.
6 Conditions for Coordinatewise Polarization
This section characterizes the information that can lead to limit coordinatewise polarization in two-dimensional state spaces. Since limit beliefs are pinned down by the priors and the identified set, the problem reduces to specifying conditions on the identified set. We now define the relevant conditions.
Definition 4 (Spanning).
A set $\Gamma\subset\Theta$ is spanning (in $\Theta$) if it obtains the lowest and highest values on each coordinate. Formally, for every coordinate $i$ there exist $\theta,\theta^{\prime}\in\Gamma$ such that $\theta_{i}=\underaccent{\bar}{\theta}_{i},\theta^{\prime}_{i}=\bar{\theta}_{i}$.
Definition 5 (Biased and balanced).
Definition 6 (Compensatory).
$\Gamma\subset\Theta$ is compensatory if there exists $\theta\in\Theta$, with $\theta\ll\bar{\theta}$ such that every state in $\Gamma$ is neither weakly below $\theta$ nor strictly above it. That is, for every $\theta^{\prime}\in\Gamma$ neither $\theta^{\prime}\leq\theta$ nor $\theta^{\prime}\gg\theta$ holds.181818 Equivalently, there exists $\hat{\theta}\in\Theta$ with $\hat{\theta}\gg\underaccent{\bar}{\theta}$ such that for every $\theta^{\prime}\in\Gamma$ neither $\theta^{\prime}\geq\hat{\theta}$ nor $\theta^{\prime}\ll\hat{\theta}$ holds. If the condition as it appears in the text is satisfied for $\theta\ll\underaccent{\bar}{\theta}$ then the version in this footnote is satisfied for $\hat{\theta}=\min\{\theta^{\prime}\in\Theta|\theta^{\prime}\gg\theta\}$.
In the characterization theorem below, the relevant condition is that a set be non-compensatory.
The theorem also requires one technical condition. We first introduce a strong form of coordinatewise dominance that rules out equalities between the marginal cdfs. Let $P^{L}$ and $P^{H}$ be two distributions whose supports are contained in $X=\{x_{1},\ldots,x_{n}\}\subset\mathbb{R}$ where $x_{1}<x_{2}<\ldots<x_{n}$. Let $F^{L}$ and $F^{H}$ denote their cdfs. Say that $P^{L}$ is strongly stochastically dominated by $P^{H}$ on $X$, denoted $P^{L}\prec^{*}_{st}P^{H}$ if $F^{L}(x_{i})>F^{H}(x_{i})$ for all $1\leq i<n$. For two distributions $P^{L},P^{H}$ whose supports are contained in $\Theta\subset\mathbb{R}^{d}$ say that $P^{L}$ is strongly coordinatewise dominated by $P^{H}$ on $\Theta$, denoted $P^{L}\prec_{cw}^{*}P^{H}$ if $P^{L}_{i}\prec^{*}_{st}P^{H}_{i}$ on $\Theta_{i}$ for all $i\in[d]$.
We now define a strong form of limit coordinatewise polarization, which requires strong dominance between the priors. Let $P^{L},P^{H}\in\Delta(\Theta)$ be priors and let $Q^{L},Q^{H}\in\Delta(\Theta)$ be the posteriors upon observing a partitional signal with realization $\Gamma$. Say that strong coordinatewise polarization occurs if
$Q^{L}\prec_{cw}P^{L}\prec^{*}_{cw}P^{H}\prec_{cw}Q^{H}$
Say that an identified set $\Gamma$ can strongly coordinatewise polarize if there exist priors for which strong coordinatewise polarization occurs for a partitional signal with realization $\Gamma$.
Theorem 5.
The conditions of Theorem 5 can be expressed geometrically. A set is spanning if it intersects every extreme row and column. A set is biased downward if it is contained in a “lower L,” while its complement is contained in the corresponding “upper L.” That is, there exists a state such that, for the Euclidean coordinate system originating at that state, $\Gamma$ covers the lower left quadrant, while $\Gamma^{C}$ covers the upper right quadrant. Similarly, a set is biased upward if it covers the upper right quadrant while its complement covers the lower left quadrant. Finally, a set is compensatory if it is contained in an off-diagonal, i.e., in the union of the upper left and lower right quadrants originating at some state.
As illustrated in Figure 1, in a $2\times 2$ state space, the unique set that satisfies all three conditions is the diagonal $\Gamma=\{(x_{1},y_{1}),(x_{2},y_{2})\}$. In general two-dimensional state spaces, any set that contains the extreme states $\underaccent{\bar}{\theta}$ and $\bar{\theta}$ is spanning and satisfies conditions (ii) and (iii), so can polarize as long its complement is spanning, i.e., as long as it does not cover any extreme row or column. In particular, in symmetric state spaces, the diagonal can always polarize while the off-diagonal cannot polarize. However, there are sets that satisfy the conditions and do not contain one or both of the extreme states. For example, in any two-dimensional product space greater than $2\times 2$ the set $\Theta\setminus\{\underaccent{\bar}{\theta},\bar{\theta}\}$ that includes all states except the two extremes satisfies all three conditions.
We now turn to the proof of Theorem 5, which requires the following result.

### v2

@@ 5 Constraints on Bayesian Polarization
Proposition 1.
Let $P,P^{\prime}\in\Delta(\Theta)$ be full support priors over a finite state space $\Theta$ and $Q,Q^{\prime}$ be the corresponding posteriors following a signal realization $x$. Assume that the likelihood function $\ell(\theta)=\Pr[X=x|\theta]$ is not constant in $\theta$. Then:
  1. 1.
    There exist two states—the minimum and maximum likelihood states—on which beliefs move in the same direction, i.e., $\left(Q(\theta)-P(\theta)\right)\left(Q^{\prime}(\theta)-P^{\prime}(\theta)\right)>0$. On all other states, beliefs may move in opposite directions.
  2. 2.
    If beliefs move in opposite directions, the divergence is consistent across states: if there exists a state $\theta\in\Theta$ for which $Q(\theta)-P(\theta)\leq 0\leq Q^{\prime}(\theta)-P^{\prime}(\theta)$, with at least one of these inequalities strict, then there does not exist a state $\theta^{\prime}$ at which $Q(\theta^{\prime})-P(\theta^{\prime})\geq 0\geq Q^{\prime}(\theta^{\prime})-P^{\prime}(\theta^{\prime})$.
Part 2 of Proposition 1 formalizes the “consistency of direction” constraint suggested by the BHK proof. It explains why the upper L creates an impossibility: stochastic polarization requires agents to reverse their direction of disagreement between the minimal and maximal states. In contrast, upper orthant polarization does not require divergence on the upper L set, so while agents’ beliefs do diverge on at least one individual state ($\bar{\theta}$), the direction of divergence is consistent across states.
Part 1 of Proposition 1 highlights that Bayesian updating otherwise imposes only mild constraints on where beliefs can diverge, allowing them to move in opposite directions on all but two states. [^22] Note that updating in opposite directions is not synonymous with polarization, as agents may update in opposite directions towards each other. But it is straightforward to construct an example where agents’ beliefs move away from each other on all but two states. This distinction helps clarify the mechanics of our earlier results. Polarization operates through two channels. The first is state-level polarization: beliefs moving in opposite directions on individual states. The second is event-level polarization: beliefs moving in the same direction at the state level, but accumulating differently on events due to differences in the magnitude of updates. In the short run, beliefs can diverge on many states but in the long-run limit they must move in the same direction on every state. Hence, one-shot polarization operates through both channels while limit polarization operates purely through the second channel.
6 Robust Polarization in Aggregate Positions
So far we have focused on polarization of multidimensional beliefs, analyzing notions that require beliefs to diverge on rich sets of events. Yet, in several applications of interest, agents’ multidimensional beliefs are mapped into one-dimensional aggregate positions, and polarization is discussed in terms of these positions. For instance, if beliefs are over the optimal policy bundle on several issue dimensions as in Example 1, with policies on each dimension ordered from most liberal to most conservative, one could also discuss polarization in terms of agents’ overall ideological positions. This aggregation of beliefs into positions could be performed by the agents themselves or by an analyst. For example, the DW-NOMINATE scaling approach uses legislators’ roll-call voting records to estimate low-dimensional ideal points. These estimates are often interpreted as ideological positions and have played a central role in studies of polarization in the U.S. Congress [26].
To formalize this type of aggregation within our model, we fix an aggregator $u:\Theta\rightarrow\mathbb{R}$ and define the agents’ prior positions to be $E_{P^{L}}[u]$ and $E_{P^{H}}[u]$. That is, an agent’s prior position is simply the expectation of $u$ under their prior. Likewise, we define their posterior positions as $E_{Q^{L}}[u]$ and $E_{Q^{H}}[u]$. We define two notions of polarization in positions upon observing a realization $x$ of the signal $X$.
Definition 8.
Given an aggregator $u:\Theta\rightarrow\mathbb{R}$, we say that $x$ leads to one-shot polarization in positions if
$E_{Q^{L}}[u]\leq E_{P^{L}}[u]\leq E_{P^{H}}[u]\leq E_{Q^{H}}[u]$ (1)
Given a set of aggregators $\mathcal{U}$, we say that $x$ leads to $\mathcal{U}$-robust one-shot polarization in positions if it leads to one-shot polarization in positions for all $u\in\mathcal{U}$, and there is a $u\in\mathcal{U}$ for which all inequalities in (1) are strict.
Limit polarization in positions and $\mathcal{U}$-robust limit polarization in positions are defined analogously, replacing beliefs after the realization $x$ with limit beliefs following a sequence of i.i.d. signals (or, equivalently, requiring that $X$ be a partitional signal).
Definition 8 considers two agents with differing beliefs whose positions are aggregated using the same aggregator $u$. This captures settings in which agents may disagree on issues but share a common convention for translating multidimensional beliefs into a one-dimensional position (e.g., two agents may disagree on many policies yet broadly agree on what constitutes a more liberal versus more conservative bundle overall). A primary motivation for considering $\mathcal{U}$-robust polarization is to remain agnostic about the exact form of the aggregator $u$. Robust polarization asks: given a class $\mathcal{U}$ of plausible aggregators, when can the same signal realization polarize aggregate positions for any $u\in\mathcal{U}$?
The following result implies a direct mapping between polarization in beliefs and robust polarization in positions (the result follows almost immediately from the definition of a generator for an integral stochastic order in Section 5). [^23] The only subtlety comes from the fact that our definition of polarization requires a strict ordering between the priors and between each prior and its posterior. This maps to the requirement that one of the functions $u\in\mathcal{U}$ leads to a strict inequality in (1). The requirement that $\mathcal{U}$ be closed under addition is mild since if $\succeq$ is generated by $\mathcal{U}$ then it is also generated by the convex cone of $\mathcal{U}$.
Proposition 2.
A signal leads to one-shot (respectively, limit) belief polarization according to an integral stochastic order $\succeq$ if and only if it leads to one-shot (respectively, limit) $\mathcal{U}$-robust polarization in positions, where $\mathcal{U}$ is a generator of $\succeq$ that is closed under addition.
Proposition 2 converts our order-based possibility/impossibility results into statements about which classes of aggregators admit robust polarization in positions. The coordinatewise order is generated by the class of additively separable increasing functions. Thus, Theorem 1 implies that for this class of aggregators, robust polarization in positions is possible both in the short run and in the limit. The upper orthant order is generated by nonnegative multiplicatively separable increasing functions. This class is not closed under addition, but by Theorem 3, robust polarization for its convex cone is possible only in the short run. In particular, there is less scope for robust polarization once the aggregator exhibits complementarities across dimensions (i.e., is multiplicatively but not additively separable). Finally, the stochastic dominance order is generated by all increasing functions. Thus, Theorem 4 implies that for aggregators that exhibit strong complementarities (i.e., are not even multiplicatively separable), robust polarization in positions is impossible after any signal realization. That is, for any signal, there exists an increasing aggregator with sufficient complementarity under which aggregate positions will not diverge.
These observations highlight a distinction between informational conditions for belief polarization and conditions on the mapping from beliefs to positions. While information about issue alignment can generate belief polarization, strong complementarities in the aggregator may prevent it from translating into polarized positions. Robust polarization in positions is facilitated by information that ties different dimensions together and aggregators that are sufficiently separable across them.
7 Conclusion
Whether polarization is rational is widely debated. To make progress, we asked instead which types of polarization are consistent with Bayesian updating from public information. For agents learning about a multidimensional state, we delineate the forms of belief divergence that are and are not possible in the short run and in the long-run limit. We find that persistent polarization of all marginal beliefs is consistent with Bayesian updating. On the other hand, we show that more demanding notions of polarization requiring divergence on combinations of issues are either impossible or short-lived. Evidence of such polarization—for example, beliefs diverging both on the event that the state is at the pointwise minimum and the event that it is at the pointwise maximum—would point to a non-Bayesian explanation.
Beyond this debate, the Bayesian model isolates the informational forces generating polarization from psychological or affective ones. Our characterization result suggests that the empirically observed rise in issue alignment may not simply be a concurrent trend, but a driver of polarization.
Appendix A Proofs
This appendix contains the proofs of all results. The proof of Theorem 2 is in Subsection A.1, along with the proofs of the lemmas used to prove this theorem.
That is, the agent’s posteriors over $\mathcal{G}$ are exactly as they would be in a learning problem where the state space is $\mathcal{G}$, the true state is $\hat{\Gamma}$, and the prior and likelihood function for each state $\Gamma$ are $P(\Gamma)$ and $\ell(\cdot;\Gamma)$, respectively. Note that for this state space it is the case that $\Gamma\neq\Gamma^{\prime}\Longrightarrow\ell(\cdot;\Gamma)\neq\ell(\cdot;\Gamma^{\prime})$ (i.e., there exists $x\in\mathcal{X}$ such that $\ell(x;\Gamma)\neq\ell(x;\Gamma^{\prime})$). Thus, by Doob’s Theorem for posterior consistency, $\lim_{t\rightarrow\infty}Q(t)(\hat{\Gamma})=1\text{ a.s.}$ [^24] The original Theorem is [11], here we apply [28] Theorem 2.4. The theorem has some technical assumptions that are easily satisfied by our finite state space. The key assumption is that the state is identifiable: $\Gamma\neq\Gamma^{\prime}\Longrightarrow\ell(\cdot;\Gamma)\neq\ell(\cdot;\Gamma^{\prime})$. This implies that for any $\theta\notin\hat{\Gamma}$, $\lim_{t\rightarrow\infty}Q(t)(\theta)=0=Q^{*}(\theta)$ a.s. Finally, for any $\theta\in\hat{\Gamma}$, following any sequence $x_{1},\ldots,x_{t}$ of signal realizations,
$\frac{Q(t)(\theta)}{Q(t)(\hat{\Gamma})}=\frac{P(\theta)\Pi_{\tau=1}^{t}\ell(x_{\tau};\hat{\Gamma})}{P(\hat{\Gamma})\Pi_{\tau=1}^{t}\ell(x_{\tau};\hat{\Gamma})}=\frac{P(\theta)}{P(\hat{\Gamma})}$
So, for any $t$, $Q(t)(\theta)=Q(t)(\hat{\Gamma})\frac{P(\theta)}{P(\hat{\Gamma})}$, which implies that $\lim_{t\rightarrow\infty}Q(t)(\theta)=\frac{P(\theta)}{P(\hat{\Gamma})}=Q^{*}(\theta)$ a.s.
Proof of Theorem 1. Let $\varepsilon\in(0,1)$, set $\Gamma=\{\underaccent{\bar}{\theta},\bar{\theta}\}$ where $\underaccent{\bar}{\theta}$ is the pointwise minimal state in $\Theta$ and $\bar{\theta}$ is the pointwise maximal state. Define
Define $P^{H}$ as a mirror image of $P^{L}$,
Note that $P^{L}_{i}\prec_{st}P^{H}_{i}$ for any $\varepsilon>0$, and that both $Q^{L}_{i}\prec_{st}P^{L}_{i}$ and $P^{H}_{i}\prec_{st}Q^{H}_{i}$ hold if and only if $\varepsilon\geq\frac{|\Theta|(n-2)}{n(|\Theta|-2)}$ (recall that we assume $n=|\Theta_{i}|\geq 2$ and $|\Theta|>2$). The right-hand side of this inequality is strictly less than $1$ since $d\geq 2$ implies $|\Theta|\geq 2n>n$. Hence, there exists $\varepsilon\in(0,1)$ such that coordinatewise polarization occurs.
Proof of Lemma 4. Let $P^{L},P^{H}\in\Delta(\Theta)$ be such that $P^{L}\preceq_{cw}P^{H}$. Let $u=\sum_{i=1}^{d}u_{i}\in\mathcal{U}$. Since $P^{L}_{i}\preceq_{st}P^{H}_{i}$ for all $i\in[d]$, we also have $E_{P^{L}_{i}}[u_{i}]\leq E_{P^{H}_{i}}[u_{i}]$ for all $i\in[d]$. So,
For the other direction, assume that $E^{L}[u]\leq E^{H}[u]$ for all $u\in\mathcal{U}$. Let $j\in[d]$ and let $u_{j}:\Theta_{j}\rightarrow\mathbb{R}$ be an increasing function. Define $u_{i}\equiv 0$ for all $i\neq j$ and $u:=\sum_{i=1}^{d}u_{i}$. Then $u(\theta)=u_{j}(\theta_{j})$ for all $\theta$, so $E^{L}[u]\leq E^{H}[u]$ implies $E_{P^{L}_{j}}[u_{j}]\leq E_{P^{H}_{j}}[u_{j}]$. Since this holds for all $j\in[d]$ and for any increasing $u_{j}:\Theta_{j}\rightarrow\mathbb{R}$ we get $P^{L}_{j}\preceq_{st}P^{H}_{j}$ for all $j\in[d]$, so $P^{L}\preceq_{cw}P^{H}$.
Proof of Theorem 3.
Impossibility of limit $\succeq_{uo}$-polarization:
Assume towards contradiction that limit $\succeq_{uo}$-polarization occurs for an identified set $\Gamma\subset\Theta$. Then $\Gamma\neq\Theta$, so for any $\theta\in\Gamma$ we have $Q^{L}(\theta)>P^{L}(\theta)$ and $Q^{H}(\theta)>P^{H}(\theta)$. The set $\{\bar{\theta}\}$ is an upper orthant in $\Theta$. If $\bar{\theta}\in\Gamma$ then $Q^{L}(\bar{\theta})>P^{L}(\bar{\theta})$, a contradiction to $Q^{L}\prec_{uo}P^{L}$. Conversely, if $\bar{\theta}\notin\Gamma$ then $Q^{H}(\bar{\theta})=0<P^{H}(\bar{\theta})$, a contradiction to $P^{H}\prec_{uo}Q^{H}$.
Possibility of one-shot $\succeq_{uo}$-polarization:
Let $n\in\mathbb{N}$ and $\varepsilon\in(0,1)$. Define priors as follows:
Let $x$ be a signal realization with likelihood function given by,
$P^{L}(U)=\frac{1}{n^{2}}+\frac{k-1}{n\left(|\Theta|-2\right)}<1-\frac{1}{n}-\frac{1}{n^{2}}+\frac{k-1}{n\left(|\Theta|-2\right)}=P^{H}(U)$
So, $P^{L}\prec_{uo}P^{H}$.
We now show that $Q^{L}\prec_{uo}P^{L}$ by proving that for sufficiently large $n$, $Q^{L}(U)<P^{L}(U)$ for any upper orthant $U\subsetneq\Theta$. By Lemma 5 below, this is equivalent to proving that for sufficiently large $n$, $E^{L}[\ell|U]<E^{L}[\ell]$ for any upper orthant $U\subsetneq\Theta$. Let $U\subsetneq\Theta$ be an upper orthant and let $k=|U|$. Then $\bar{\theta}\in U$ and $\underaccent{\bar}{\theta}\notin U$ imply,
On the other hand, as $n\rightarrow\infty$
Since $E^{H}[\ell]$ is a weighted average of $E^{H}[\ell|U]$ and $E^{H}[\ell|\Theta\setminus U]$ to prove that $E^{H}[\ell|U]>E^{H}[\ell]$ it suffices to prove that $E^{H}[\ell|U]>E^{H}[\ell|\Theta\setminus U]$. This holds for sufficiently large $n$, since as $n\rightarrow\infty$,
For the limit, we used the fact that $|U|=k<|\Theta|-1$. To see this, first note that for any upper orthant $U\subsetneq\Theta$ it must be the case that $\underaccent{\bar}{\theta}\notin U$. So, if $|U|\geq|\Theta|-1$, then $U=\Theta\setminus\{\underaccent{\bar}{\theta}\}$. Let $a\in\mathbb{R}^{d}$ be the origin of the upper orthant $U$, i.e., the point for which $U=\{\theta\in\Theta:\theta_{1}>a_{1},\theta_{2}>a_{2},\ldots,\theta_{d}>a_{d}\}$. Let $i\in[d]$. Since $U=\Theta\setminus\{\underaccent{\bar}{\theta}\}$ there exists a state $\theta\in U$ with $\theta_{i}=\underaccent{\bar}{\theta}_{i}$. Thus $a_{i}\leq\theta_{i}$. Since this holds for all $i\in[d]$ we have $a\leq\underaccent{\bar}{\theta}$, which implies $\underaccent{\bar}{\theta}\in U$, a contradiction.
Proof of Lemma 5. Let $A\subset\Theta$ be a nonempty set. Since $Q$ is obtained by Bayesian updating with prior $P$ and likelihood $\ell$, we have,
Proof of Theorem 4. Assume towards contradiction that there exists a signal realization $x$ that leads to one-shot $\succeq_{st}$-polarization. We will prove by induction on $n$ that for every upper set $U\subset\Theta$ of size $n$: $P^{L}(U)=Q^{L}(U)$ and $P^{H}(U)=Q^{H}(U)$. This contradicts the definition of $\succeq_{st}$-polarization, which requires the existence of upper sets $U$ and $U^{\prime}$ such that $P^{L}(U)>Q^{L}(U)$ and $P^{H}(U^{\prime})<Q^{H}(U^{\prime})$.
Base case: Since $\{\bar{\theta}\}$ and $\Theta\setminus\{\underaccent{\bar}{\theta}\}$ are upper sets, the assumption $Q^{L}\prec_{st}P^{L}$ implies that $Q^{L}(\bar{\theta})\leq P^{L}(\bar{\theta})$ and $Q^{L}(\underaccent{\bar}{\theta})\geq P^{L}(\underaccent{\bar}{\theta})$. Similarly, the assumption $Q^{H}\succ_{st}P^{H}$ implies $Q^{H}(\bar{\theta})\geq P^{H}(\bar{\theta})$ and $Q^{H}(\underaccent{\bar}{\theta})\leq P^{H}(\underaccent{\bar}{\theta})$. By Proposition 1, none of these inequalities can be strict, so $Q^{L}(\bar{\theta})=P^{L}(\bar{\theta})$ and $P^{H}(\bar{\theta})=Q^{H}(\bar{\theta})$. Since $\{\bar{\theta}\}$ is the unique upper set of size $1$ this completes the base case. [^25] This theorem also holds without the assumption that $\Theta$ is a product space, in which case it may not have a unique maximum. The only change is in the base case, where we would let $\bar{\theta}$ be some maximal element of $\Theta$ and $\underaccent{\bar}{\theta}$ be some minimal element of $\Theta$.
Inductive step: Let $1<n\leq|\Theta|$, and assume that the inductive hypothesis holds for all $1\leq n^{\prime}<n$. Consider an upper set $U$ of size $n$. Let $\theta_{U}$ be a minimal element of $U$ and define $A=U\setminus\{\theta_{U}\}$. Then $A$ is an upper set of size n1n-1. The assumption that $Q^{L}\prec_{st}P^{L}$ implies $P^{L}(A)+P^{L}(\theta_{U})=P^{L}(U)\geq Q^{L}(U)=Q^{L}(A)+Q^{L}(\theta_{U})$. The inductive hypothesis implies $P^{L}(A)=Q^{L}(A)$, which implies $P^{L}(\theta_{U})\geq Q^{L}(\theta_{U})$. By an analogous argument $P^{H}(\theta_{U})\leq Q^{H}(\theta_{U})$. So, as in the base case, Proposition 1 implies $P^{L}(\theta_{U})=Q^{L}(\theta_{U})$ and $P^{H}(\theta_{U})=Q^{H}(\theta_{U})$. Hence, $P^{L}(U)=Q^{L}(U)$ and $P^{H}(U)=Q^{H}(U)$, which completes the induction argument.
Proof of Proposition 1. Let $\theta^{-}\in\arg\min(\ell)$ and $\theta^{+}\in\arg\max(\ell)$ be states attaining the minimum and maximum likelihood values. Let $E_{P}[\ell],E_{P^{\prime}}[\ell]$ denote the expectations of $\ell$ under $P$ and $P^{\prime}$, respectively. By Lemma 5, for every $\theta\in\Theta$ we have $Q(\theta)>P(\theta)$ if and only if $\ell(\theta)>E_{P}[\ell]$, and similarly for $Q^{\prime},P^{\prime}$. Since $\ell$ is not constant, we have $\ell(\theta^{-})<\ell(\theta^{+})$. Because $P$ and $P^{\prime}$ have full support, it follows that $\ell(\theta^{-})<\min\{E_{P}[\ell],E_{P^{\prime}}[\ell]\}$ and $\ell(\theta^{+})>\max\{E_{P}[\ell],E_{P^{\prime}}[\ell]\}$. Thus, for each $\theta\in\{\theta^{-},\theta^{+}\}$ we have $(Q(\theta)-P(\theta))(Q^{\prime}(\theta)-P^{\prime}(\theta))>0$.
Finally, assume that there exists $\theta\in\Theta$ for which $Q(\theta)-P(\theta)<0\leq Q^{\prime}(\theta)-P^{\prime}(\theta)$ (here the first inequality is the strict one w.l.o.g.). Then, $E_{P}[\ell]>\ell(\theta)\geq E_{P^{\prime}}[\ell]$. So, there does not exist $\theta^{\prime}\in\Theta$ at which $Q(\theta^{\prime})-P(\theta^{\prime})\geq 0\geq Q^{\prime}(\theta^{\prime})-P^{\prime}(\theta^{\prime})$, as that would imply $E_{P}[\ell]\leq E_{P^{\prime}}[\ell]$, a contradiction.
Proof of Proposition 2. Let $\succeq$ be an integral stochastic order over $\Delta(\Theta)$ generated by a class $\mathcal{U}$ of functions $u:\Theta\rightarrow\mathbb{R}$. We prove the one-shot case; the limit case is analogous. Assume that one-shot $\succeq$-polarization is possible. Then there exist priors and a signal realization such that $Q^{L}\prec P^{L}\prec P^{H}\prec Q^{H}$. By definition of a generator, this implies that (1) holds for any $u\in\mathcal{U}$. It remains to prove that there exists $u\in\mathcal{U}$ for which all inequalities in (1) are strict. Since $Q^{L}\prec P^{L}$, there exists $u_{1}\in\mathcal{U}$ such that $E_{Q^{L}}[u_{1}]<E_{P^{L}}[u_{1}]$ (otherwise $E_{Q^{L}}[u]\geq E_{P^{L}}[u]$ for all $u\in\mathcal{U}$, implying $Q^{L}\succeq P^{L}$, a contradiction). Similarly, there exist $u_{2},u_{3}$ in $\mathcal{U}$ such that $E_{P^{L}}[u_{2}]<E_{P^{H}}[u_{2}]$ and $E_{P^{H}}[u_{3}]<E_{Q^{H}}[u_{3}]$. Let $u=u_{1}+u_{2}+u_{3}$. Then, $u\in\mathcal{U}$ (since $\mathcal{U}$ is closed under addition) and is such that all inequalities in (1) are strict. For the converse, assume that one-shot $\mathcal{U}$-robust polarization in positions is possible. Then, similar to above, since there exists a signal for which (1) holds for any $u\in\mathcal{U}$, we have $Q^{L}\preceq P^{L}\preceq P^{H}\preceq Q^{H}$. Since there exists a $u\in\mathcal{U}$ for which (1) holds strictly, it must be the case that $Q^{L}\prec P^{L}\prec P^{H}\prec Q^{H}$.
A.1 Proof of Theorem 2
Proof of Theorem 2. Since $\Theta\subset\mathbb{R}^{2}$ we can assume $\Theta=\{x_{1},\ldots,x_{m}\}\times\{y_{1},\ldots,y_{n}\}$ for $m,n\geq 2$, where $x_{1}<\ldots<x_{m}$ and $y_{1}<\ldots<y_{n}$. Throughout this proof, let $F^{L}_{i},F^{H}_{i}$ denote the cdfs of the marginal distributions of $P^{L}$ and $P^{H}$, respectively on the $i$ coordinate. Let $F^{L}_{i}|_{\Gamma},F^{L}_{i}|_{\Gamma}^{C}$ denote the cdfs of the marginal distributions of $P^{L}|_{\Gamma}$ and $P^{L}|_{\Gamma^{C}}$, and similarly define $F^{H}_{i}|_{\Gamma},F^{H}_{i}|_{\Gamma^{C}}$.
The proof uses Lemma 2 to translate polarization into a condition on priors, as well as the following two results on strong stochastic dominance and strong coordinatewise dominance. Proofs of all lemmas stated in the course of the theorem proof appear at the end of this subsection.
@@ A.1 Proof of Theorem 2
Let $P^{L}$ and $P^{H}$ be two univariate distributions with a common finite support $X\subset\mathbb{R}$. Then $P^{L}\prec^{*}_{st}P^{H}$ implies that $E^{L}[u]<E^{H}[u]$ for any function $u:X\rightarrow\mathbb{R}$ that is weakly increasing and not constant.